# FastSLM: Hierarchical Frame Q-Former for Effective Speech Modality Adaptation

### v1

@@ Appendix F Generation and Decoding Configuration
Appendix G Model and Training Parameters
@@
Appendix H Details of ASR Benchmark Results
@@
OpenASR AMI WER $12.8$ $15.2$ $11.7$ $16.0$ $16.3$ $21.6$
Earnings22 WER $10.5$ $14.1$ $10.2$ $11.3$ $10.7$ $13.1$
GigaSpeech WER $11.2$ $10.3$ $9.78$ $10.0$ $10.2$ $10.7$
SpgiSpeech WER $2.52$ $3.00$ $3.13$ $2.01$ $2.37$ $3.82$
TEDLIUM WER $3.84$ $4.05$ $2.90$ $3.91$ $3.68$ $3.01$
@@
Appendix I Mathematical Justification for Speech Token Ratio
@@
Appendix K Use of AI Writing Assistance
LLMs were used solely to refine the language and clarity of this manuscript. All technical content was written by the humans, and all AI-assisted edits were reviewed and approved by them.

### v2

@@ Appendix G Generation and Decoding Configuration
Appendix H Model and Training Parameters
@@
Appendix I Details of ASR Benchmark Results
@@
OpenASR AMI WER $10.8$ $15.2$ $11.7$ $16.0$ $16.3$ $21.6$
Earnings22 WER $10.7$ $14.1$ $10.2$ $11.3$ $10.7$ $13.1$
GigaSpeech WER $10.7$ $10.3$ $9.78$ $10.0$ $10.2$ $10.7$
SpgiSpeech WER $2.33$ $3.00$ $3.13$ $2.01$ $2.37$ $3.82$
TEDLIUM WER $3.97$ $4.05$ $2.90$ $3.91$ $3.68$ $3.01$
@@
Appendix J Mathematical Justification for Speech Token Ratio
@@
Appendix L Use of AI Writing Assistance
LLMs were used solely to improve the language and clarity of this manuscript. All technical content was authored by the human authors, and all AI-assisted edits were reviewed and approved by the authors.